OpenStudy (anonymous):

can you help me. [7.01] Choose the polynomial written in standard form. x2y2 + 4x4y + 10x6 x4y2 + 4x3y5 + 10x x7y2 + 4x3y + 10x3 x6y2 + 4x3y8 + 10x2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

maybe this one x4y2 + 4x3y5 + 10x

OpenStudy (skullpatrol):

Why?

OpenStudy (anonymous):

@skullpatrol can you explain this? Even if @Felicia123 gets the right answer please?

OpenStudy (skullpatrol):

Do you know the definition for the standard form of a polynomial?

OpenStudy (anonymous):

in descending order from highest to lowest so it is the third one.

OpenStudy (anonymous):

you have to put the exponents in order. :)

OpenStudy (anonymous):

you sure its 3

OpenStudy (anonymous):

Yep positive

OpenStudy (anonymous):

How do you determine which term has the highest exponents when there are multiple variables?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

Neither do I. I was asking @Maxicat or @skullpatrol - These questions have never made any sense to me.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

you make sure the powers are in decreasing order you have to add the exponents of the variable then decide. i know its confusing

OpenStudy (anonymous):

yes it is lol

OpenStudy (anonymous):

I promise the answer is the third one

OpenStudy (anonymous):

Ok, so you add the exponents of all variables in a term to get an index number (not sure what to call it) and list the terms in descending index value.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What is the point of writing an expression or equation in this form?

OpenStudy (skullpatrol):

They are called "exponents"

OpenStudy (anonymous):

refers to the constant term in the order

OpenStudy (anonymous):

Zero refers, sorry, screen keeps jumping all over lol

OpenStudy (anonymous):

so why is it 3 one

OpenStudy (anonymous):

\( x^7y^2 + 4x^3y + 10x^3\) First term index = 7 + 2 = 9 Second term index = 3 + 1 = 4 Third term index = 3 It is in descending order according to the index value.

OpenStudy (anonymous):

\(x^2y^2 + 4x^4y + 10x^6\) First term index = 2 + 2 = 4 Second term index = 4 + 1 = 5 Third term index = 6 Not in descending order.

OpenStudy (anonymous):

oh i got it now thanks